OpenStudy (haltiamali):

9x^4-27x^2+20=0 help please? gold medal to the person that works it out for me!

OpenStudy (aum):

Try the AC method to factor the left hand side. 9x^4-27x^2+20=0 a = 9, b = -27, c = 20 a * c = 9 * 20 = 180 See if you can find factors of 180 that add to -27. Obviously both factors have to be negative here.

OpenStudy (aum):

Or split -27 into two parts that add to -27 and multiplies to +180.

OpenStudy (haltiamali):

How do I split it?

OpenStudy (aum):

-15 and -12

OpenStudy (haltiamali):

So then how would the equation look then? 9x^4-15-12+20=0?

OpenStudy (haltiamali):

Oops 9x^4-15x-12x+20

OpenStudy (aum):

9x^4-27x^2+20=0 9x^4-12x^2-15x^2+20=0 3x^2(3x^2-4) -5(3x^2-4) = 0 (3x^2-5)(3x^2-4) = 0 3x^2 - 5 = 0 or 3x^2 - 4 = 0 3x^2 = 5 or 3x^2 = 4 x^2 = 5/3 or x^2 = 4/3 \(\large x = -\sqrt{\frac 53}; ~~~~x = \sqrt{\frac 53}; ~~~~ x = -\sqrt{\frac 43}=-2\sqrt{\frac 13}; ~~~~x = \sqrt{\frac 43} =2\sqrt{\frac 13}\)

OpenStudy (haltiamali):

thank you!